OpenStudy (anonymous):

By method of partial fractions: Y = s/[(s^2+1)(s^2+2s+2)]

OpenStudy (amistre64):

well, we start by factoring the denominator to see what we have to use

OpenStudy (amistre64):

this looks to be factored well enough; so split it into 2 fractions: \[\frac{s}{(s^2+1)(s^2+2s+2)}=\frac{Ax+B}{s^2+1}+\frac{Cx+D}{s^2+s2s +2}\]

OpenStudy (amistre64):

it looks better without the typo ;)